If a,b,c in A.P. prove that. a×a(b+c),b×b(c+a),c×c(a+b) are in A.P.

Grade:11

1 Answers

Arun
6 years ago

Given, a,b and c are in A.P.

a2(b + c), b2(c + a), c2(a + b) will be in A.P.

if b2(c + a) – a2(b + c) = c2(a + b) – b2(c + a)

i.e., if c(b2 – a2) + ab(b – a) = a(c2 – b2) + bc(c– b)

i.e., if (b – a)(ab + bc + ca) = (c – b)(ab + bc + ca)

i.e., if b – a = – b

i.e., if 2b = a + c

i.e., if abare in A.P.

Thus, if abc are in A.P. then a2(b + c), b2(c + a), c2(a + b) will also be in A.P.